Given that triangle ABC is similar to triangle FED,where AB = 77,EF = 21, BC = 11x + 11, and DE = 30.Find x.
Answer by Theo(13342)

if the triangles are similar, then the sides are proportional.

this means that AB / EF = BC / DE

AB= 77
EF = 21
BC = 11x + 11
DE = 30

ratio becomes:

77 / 21 = (11x + 11) / 30

multiply both sides of this equation by 30 and you get:

(77 * 30) / 21 = 11x + 11

simplify this to get:

110 = 11x + 11

subtract 11 from both sides of this equation to get:

99 = 11x

divide both sides of this equation by 11 to get:

9 = x

that's your answer.

11x + 11 is equal to 99 + 11 is equal to 110.

the ratio of:

77 / 21 = (11x + 11) / 30 becomes:

77 / 21 = 110 / 30

cross multiply to get 77 * 30 = 21 * 110

simplify to get 2310 = 2310, confirming the solution of x = 9 is good.